# thanks: Jointly supervised this workthanks: Jointly supervised this work

### v2

@@ Abstract
Figure 1: Schematic of a non-Hermitian lateral TIJJ: A ‘short’ TIJJ formed by placing two bulk $s$-wave singlet superconductors on the surface of a 3DTI. A top gate and a ‘lossy’ metallic lead ($\mathcal{N}$) are attached at the junction ($x=0$), where the gate modifies the potential barrier via an applied voltage $V_{g}$ and the lead introduces dissipation due to coupling to an electron reservoir.
In this work, we address the central question: how does dissipation caused by coupling a ‘lossy’ metallic lead [cayao2023non, Ohnmacht2024] at the junction, affect the quantum transport properties of TIJJs? To explore this, we model the dissipative effects of the lead by an effective non-Hermitian Hamiltonian [Harrington2022, Li2024] and investigate transport in a non-Hermitian lateral Josephson junction on the surface of a 3DTI (NH-TIJJs), as illustrated schematically in Fig. 1. Here, proximity-induced superconductivity intertwines with non-Hermitian effects, offering a unique platform to study the interplay between dissipation and topological superconductivity.
@@
Figure 2: Characteristics of the NH-TIJJ for a complex barrier ($Z_{1}\neq 0$, $Z_{2}\neq 0$) in the tunneling limit ($\Gamma_{0}\ll\Delta$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the NH-TIJJ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$ taking the imaginary barrier parameter $Z_{2}=0.2$. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the asymmetric gap-exit points in the spectrum. (c) The zero temperature angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. The current is normalized by $I_{0}=eN\Gamma_{0}/2\hbar$ and we have chosen the real barrier parameter $Z_{1}=0.4$ and tunneling parameter $\gamma=0.3$.
@@
Figure 3: (a) Variation of the imaginary part of the Andreev spectrum as a function of the real barrier parameter $Z_{1}$ and phase difference $\phi$. (b) The total current plotted as a function of $\phi$ for different values of the barrier parameter $Z_{1}$. The imaginary barrier parameter is fixed at $Z_{2}=0.2$ in both cases.
Figure 4: Characteristics of the ordinary NH-$J$ for a general complex barrier ($Z_{1}\neq 0,Z_{2}\neq 0$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the 2D NH-$J$ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$ taking the imaginary barrier parameter $Z_{2}=0.2$. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the symmetric gap-exit points in the spectrum. (c) The angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence at zero temperature. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. Other parameters are the same as given in Fig. 2.
@@
$I_{\theta}(\phi)=\frac{e\Delta}{\hbar}k_{B}T\sum_{\omega_{n}}\frac{1}{\Omega_{n}}[r_{he}(\phi,\omega_{n})-r_{eh}(\phi,\omega_{n})]\;,$ (8)
where analytic continuation for the energy is done as $E\to i\omega_{n}$ in terms of the Matsubara frequencies $\omega_{n}=(2n+1)\pi/k_{B}T$ with $n$ being an integer, the parameter $\Omega_{n}=\sqrt{\omega^{2}_{n}+\Delta^{2}}$, $r_{he}$ is the Andreev coefficient for an incoming electron-like quasiparticle being reflected as a hole-like quasiparticle and $r_{eh}$ is the corresponding coefficient for the reverse process for a hole-like quasiparticle. In the zero-temperature limit $T\to 0$, the summation over Matsubara frequencies can be replaced by an integration over real energy as $k_{B}T\sum_{\omega_{n}}\to\frac{1}{2\pi}\int d\omega$ on performing the analytic continuation $i\omega_{n}\to\omega+i0^{+}$. We note that the existing supercurrent formulae for NH systems used to evaluate the current in systems with exceptional points [Pino2025, Shen2024] incur a divergence of the angle-resolved current at the GEPs, therefore to resolve these divergences we use the FT formula and limit our analysis within the subgap approximation ($|\omega|<\Delta$). The variation of the angle-resolved current $I_{\theta}(\phi)$ for different $\theta$ at zero temperature within the subgap approximation is shown in Fig. 2(c). The current is $2\pi$-periodic and exhibits an ‘anomalous’ behavior where the supercurrent is non-zero at $\phi=0,2\pi$ for all angles of incidence. For normal incidence, the current changes sign smoothly and does not exhibit the abrupt discontinuity at the zero-energy crossing as in the Hermitian case, which highlights the contribution of the imaginary part of the spectrum in reshaping and smoothing the CPRs. For oblique incidence, in addition to the mentioned characteristics, the critical current decreases with increasing $\theta$ indicating that the effect of non-Hermiticity is more effective with increasing obliqueness.
@@
Comparison with ‘ordinary’ NH-JJs: We consider an ‘ordinary’ lateral NH-JJ, where the 3DTI in Fig. 1 is replaced by an ‘ordinary’ metal, to highlight the special features of the NH-TIJJ. Proceeding in a similar way, in the ‘short’ junction limit within the Andreev approximation, the ABS energies are given by (see Appendix C for details),
@@
Figure 5: Comparison of the CPR computed from the Furusaki-Tsukada (FT) formula (Eq. (8)) and the Andreev-bound-state (ABS) formula (Eq. (11)) for: (a) planar ordinary NH-$J$. (b) strong-proximity NH-TIJJ. The barrier parameters are fixed at $Z_{1}=0.4,Z_{2}=0.2$ in both cases.
Comparison of CPR from ABS formula: Given the inability of the alternative supercurrent formulae for NH systems with EPs to resolve the divergence of the supercurrent at the GEPs in our system as mentioned earlier, we calculate the supercurrent from the ABS formula derived from the polar Green’s function [Capecelatro2025] and compare with that computed from the FT formula. For this purpose, we numerically compute the angle-resolved Josephson current at zero temperature as [Capecelatro2025]:
@@
after decomposing the Andreev spectrum as $E_{b}=\epsilon-i\lambda$ with $\epsilon$ and $\lambda$ being the real and imaginary part of the spectrum respectively. The resulting CPR is plotted in Fig. 5 for both the NH-TIJJ and NH-JJ. Within the subgap approximation, the CPR computed from the polar ABS formula exhibits intervals along $\phi$, where the supercurent vanishes, called Josephson gaps. The appearance of Josephson gaps outside the GEPs is very similar to the results shown in Ref. [Li2024]. However, notice that there is a sizable discrepancy between the two approaches, more prominently around the GEPs, there are no Josephson gaps in the CPR computed using the FT formula. This mismatch arises because the ABS formula takes into account only the contribution from the bound-state spectrum which comes from the poles of the Green’s function, thereby resulting in the absence of supercurrent outside the GEPs as there are no bound states. However, in non-Hermitian systems, there is another contribution stemming from the branch-cut part of the Green’s function [Capecelatro2025]. This latter contribution is not captured by the polar ABS formula but is included in the FT formula since the current is computed by taking into account all the reflection processes, thus giving a finite supercurrent outside the GEPs.

### v3

@@ Abstract
Figure 1: Schematic of a non-Hermitian planar TIJJ: A ‘short’ TIJJ formed by placing two bulk $s$-wave singlet superconductors on the surface of a 3DTI. A top gate and a ‘lossy’ metallic lead ($\mathcal{N}$) are attached at the junction ($x=0$), where the gate modifies the potential barrier via an applied voltage $V_{g}$ and the lead introduces dissipation due to coupling to an electron reservoir.
In this work, we address the central question: how does dissipation caused by coupling a ‘lossy’ metallic lead [cayao2023non, Ohnmacht2024] at the junction, affect the quantum transport properties of TIJJs? To explore this, we model the dissipative effects of the lead by an effective non-Hermitian Hamiltonian [Harrington2022, Li2024] and investigate transport in a non-Hermitian planar Josephson junction on the surface of a 3DTI (NH-TIJJs), as illustrated schematically in Fig. 1. Here, proximity-induced superconductivity intertwines with non-Hermitian effects, offering a unique platform to study the interplay between dissipation and topological superconductivity.
@@
Figure 2: Characteristics of the NH-TIJJ for a complex barrier ($Z_{1}\neq 0$, $Z_{2}\neq 0$) in the tunneling limit ($\Gamma_{0}\ll\Delta$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the NH-TIJJ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$ taking the imaginary barrier parameter $Z_{2}=0.2$. The $\phi_{1}$ and $\phi_{2}$ denote the GEPs, beyond which the real part of the spectrum exits the gap and the imaginary part of the spectrum vanishes. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the asymmetric gap-exit points in the spectrum. (c) The zero temperature angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. We have chosen the real barrier parameter $Z_{1}=0.4$, tunneling parameter $\gamma=0.3$ and $I_{0}=eN\Gamma_{0}/(2\hbar)$ with $N$ being the number of open channels.
@@
Figure 3: Characteristics of the NH-TIJJ for fixed $Z_{2}=0.2$ in the tunneling limit: Variation of the (a) imaginary part of the Andreev spectrum as a function of $Z_{1}$ and $\phi$, and (b) total current as a function of $\phi$ for different values of $Z_{1}$.
Figure 4: Characteristics of the ordinary NH-$J$ for a complex barrier ($Z_{1}\neq 0,Z_{2}\neq 0$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the 2D NH-$J$ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$ taking the imaginary barrier parameter $Z_{2}=0.2$. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the symmetric gap-exit points in the spectrum. (c) The angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence at zero temperature and $Z_{2}=0.2$. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. Here, $I^{\prime}_{0}=e\Delta/\hbar$ and $Z_{1}=0.4$.
@@
$I_{\theta}(\phi)=\frac{e}{\hbar}k_{B}T\sum_{\omega_{n}}\frac{\tilde{\Delta}(\omega_{n})}{\Omega_{n}}[r_{he}(\phi,\omega_{n})-r_{eh}(\phi,\omega_{n})]\;,$ (8)
where analytic continuation for the energy is done as $E\to i\omega_{n}$ in terms of the Matsubara frequencies $\omega_{n}=(2n+1)\pi/k_{B}T$ with $n$ being an integer, the parameter $\Omega_{n}=\sqrt{\omega^{2}_{n}+\Delta^{2}}$, $r_{he}$ is the Andreev coefficient for an incoming electron-like quasiparticle being reflected as a hole-like quasiparticle and $r_{eh}$ is the corresponding coefficient for the reverse process for a hole-like quasiparticle. In the zero-temperature limit $T\to 0$, the summation over Matsubara frequencies can be replaced by an integration over real energy as $k_{B}T\sum_{\omega_{n}}\to\frac{1}{2\pi}\int d\omega$ on performing the analytic continuation $i\omega_{n}\to\omega+i0^{+}$. We note that the existing supercurrent formulae for NH systems used to evaluate the current in systems with EPs [Pino2025, Shen2024] incur a divergence of the angle-resolved current at the GEPs, therefore to resolve these divergences we use the FT formula and limit our analysis within the subgap approximation ($|\omega|<\Delta$). The variation of the angle-resolved current $I_{\theta}(\phi)$ for different $\theta$ at zero temperature within the subgap approximation is shown in Fig. 2(c). The current is $2\pi$-periodic and exhibits an ‘anomalous’ behavior where the supercurrent is non-zero at $\phi=0,2\pi$ for all angles of incidence. For normal incidence, the current changes sign smoothly and does not exhibit the abrupt discontinuity at the zero-energy crossing as in the Hermitian case, which highlights the contribution of the imaginary part of the spectrum in reshaping and smoothing the CPRs. For oblique incidence, in addition to the mentioned characteristics, the critical current decreases with increasing $\theta$ indicating that the effect of non-Hermiticity is more effective with increasing obliqueness.
@@
Comparison with ‘ordinary’ NH-JJs: We consider an ‘ordinary’ planar NH-JJ, where the 3DTI in Fig. 1 is replaced by an ‘ordinary’ metal, to highlight the special features of the NH-TIJJ. Proceeding in a similar way, in the ‘short’ junction limit within the Andreev approximation, the ABS energies are given by (see Appendix C for details),
@@
Figure 5: Comparison of the CPR obtained from the FT formula (Eq. (8)) and the ABS formula (Eq. (11)) for a complex barrier ($Z_{1}=0.4$ and $Z_{2}=0.2$): (a) Planar ordinary NH-$J$, and (b) NH-TIJJ in the strong-proximity regime.
Comparison with CPR from ABS formula: The alternative supercurrent formulae for NH systems [Pino2025, Shen2024] with EPs are not able to resolve the divergence of the supercurrent at the GEPs in our system as mentioned earlier. Therefore we calculate the supercurrent from an ABS formula derived from the polar Green’s function [Capecelatro2025] focusing only on the subgap contribution and compare with the CPR computed from the FT formula. The angle-resolved Josephson current at zero temperature from this ABS formula is given by [Capecelatro2025]:
@@
after decomposing the Andreev spectrum as $E_{b}=\epsilon-i\lambda$ with $\epsilon$ and $\lambda$ being the real and imaginary part of the spectrum respectively. The resulting CPR is plotted in Fig. 5 for both the NH-JJ and the NH-TIJJ in the strong-proximity regime. Within this subgap approximation, the ABS formula predicts ‘Josephson gaps’, intervals of vanishing supercurrent [Li2024]. However, a significant discrepancy exists between these two formulas, particularly near the GEPs where the FT result shows no such gaps as shown in Fig. 5. This mismatch arises because the ABS formula captures only the contribution from the bound-state spectrum (Green’s function poles), yielding zero current where bound states are absent (beyond GEPs). In non-Hermitian systems, however, a critical contribution stems from the Green’s function branch cuts [Capecelatro2025]. Unlike the ABS formula, the FT approach incorporates all the reflection processes, thereby recovering a finite supercurrent outside the GEPs as well.